OpenStudy (anonymous):

Suppose you have 30 CDs. You know that you have 11 more CDs than your friend. Write and solve an equation to find the number of CDs your friend has.

OpenStudy (anonymous):

let the no of CD yr friend has = x then you have x +11 cds x+ 11 = 30

OpenStudy (anonymous):

x = 30-11 x = 19 so yr frnd has 19 CDs

OpenStudy (anonymous):

can u pls become my fan???? i will be obliged.... o( :))

OpenStudy (anonymous):

thank YOU !!!!

OpenStudy (anonymous):

:)